OpenStudy (anonymous):

find the domain of the function g(x)= 7/8-7x

OpenStudy (nenadmatematika):

I assume you mean 7/(8-7x)?

OpenStudy (nenadmatematika):

if so, the solution is: \[x \in(-\infty,8/7)\cup(8/7,\infty)\]

OpenStudy (anonymous):

it is a fraction 7 over 8-7x

OpenStudy (nenadmatematika):

then you have a correct solution above :D

OpenStudy (anonymous):

multiple choice a. real number and x =(with a slash through it) 7 b. x=(with slash) 8/7 c. x \[\ge\] 8/7 d. x=(with slash)0